# LLMatic: Neural Architecture Search via Large Language Models and Quality Diversity Optimization

### v7

@@ 1. Introduction
One could argue that search by evolutionary computation or reinforcement learning is quite mindless and wasteful, given how many architectures need to be tested and how uninformed the changes that lead to each new architecture are. Is there some way we can inform the search by exploiting stored knowledge about how to design neural networks? This paper explores the idea that we can do exactly this using code-generating large language models (LLMs). More precisely, we propose using LLM with evolutionary algorithm to generate new architectures that have high network architectural diversity and state-of-the-art performance.
@@ 2. Related Work
To avoid having the designer constrain the space of possible architectures prior to search, we turn to code-generating Large Language Models (LLMs)—large models trained auto-regressively on massive datasets of code (e.g. public repositories hosted on Github). These LLMs are based on the transformer architecture (Vaswani et al., 2017) and have obtained state-of-the-art performance in natural language modelling (Radford et al., 2019; Brown et al., 2020; Adelani et al., 2022; Nasir and Mchechesi, 2022). They have also been successfully used in specific applications, such as for video game level design (Todd et al., 2023; Nasir and Togelius, 2023; Chen et al., 2021) or code generation (Shinn et al., 2023).
@@ 3. Approach
1  Initialize network and prompt archives. while Number of generations $<$ maximum generations
@@
4                   Mutate the initial network with a random prompt. Get the network and prompt individuals and add them to the respective archives.
@@
6                   randomly choose mutation or crossover as the evolutionary operator. if evolutionary operator ==absent$=$ = mutation then
7                         Select network and prompt. Mutate the selected network with the selected prompt. Train or query the generated network. Get the network and prompt individuals and store them.
@@
9                         Selection of networks for crossover. Perform crossover on selected networks with a fixed prompt. Train or query the generated network. Store the generated network individual.
@@
15      Evaluate all networks to find the test accuracies. Mutate the temperature for the next generation. Add the batch network individuals and the corresponding prompt individuals in the respective archives.
@@
LLMatic begins its search with a very basic neural network, inspired by the work of  (Stanley and Miikkulainen, 2002) which suggests that neuroevolution tends to perform better when starting with a small network. In LLMatic, we use a novel dual-archive cooperative QD optimization approach, in which two separate archives are used to store complementary components that can be combined to solve a given task. The first archive stores neural networks, where the width-to-depth ratio and Floating Point Operations per Second (FLOPS) of a network are the behavioural descriptors. The width-to-depth ratio is a division of the width and the depth of the network. To specify the width, we use the maximum of the output features of all layers, while depth is simply the number of layers. Note that we choose FLOPS instead of parameter count because FLOPS correlates better with actual time spent training a network (Ayala et al., 2017). We call this archive the “network archive”. The fitness function for the networks in this archive is defined as the test accuracy of the network after training. The second archive, called the “prompt archive”, contains the prompt and temperature used for generating code, which are also the behavioural descriptors. Temperature of an LLM is a parameter that controls the creativity of the LLM. Lower temperature means the LLM will not diverge from the answer, higher means it will be more creative. The selection of prompt and temperature depends on a curiosity score (Cully and Demiris, 2017), governed by whether the generated network was added to the network archive. The fitness of prompt individuals depends on whether the network was better than the previous generation’s best score. Figure 1 illustrates the flowchart of the approach LLMatic uses.
In the first generation, a simple neural network with one convolutional and one fully connected layer initiates the evolution [line 1 of algorithm 1]. A prompt is selected at random to generate an initial batch of networks [line 5-6 of algorithm 1]. These networks are evaluated and an attempt is made to add them to the network archive as a random initialization for MAP-Elites. Concurrently, we mutate the temperature based on the fitness of the network, increasing it if the fitness increases and vice versa [line 21-23 of algorithm 1]. By increasing the temperature, we want the LLM to explore as it is performing better than before. By decreasing the temperature, we want the LLM to exploit and try to achieve better fitness than before. Once we calculate the fitness of the prompt individual, we add the score to a collective prompt fitness score, after which we try to populate the prompt archive. The collective prompt fitness score determines the overall fitness of each individual in the prompt archive as it gives each prompt a fitness score.
Once either of the archives reaches a specified capacity, we introduce neural network training and evolutionary operators in the process [line 7-20 of algorithm 1]. With a certain probability at each generation, a decision is made on whether to perform crossover or mutation to produce $N$ new offspring. If the crossover operator is chosen, we select $N$ random network individuals, locate their closest networks in the archive, and carry out a crossover operation instructed by a prompt [line 15-16 of algorithm 1]. No individual is added to the prompt archive when a crossover is performed. If the mutation operation is selected, we pick the most curious prompt individual and a random network individual. For exploration, we also select random prompts. In both cases, each network is trained for a certain number of epochs and an attempt is made to add the network to the archive. Likewise, a prompt individual is added as previously described. This process continues for a pre-determined number of generations. Algorithm 1 shows the complete search process of LLMatic in pseudocode. Refer to supplementary material for pseudocode on mutation operators, crossover operators, temperature mutation and addition to archives.
@@ 4.1. Setting up LLMatic
Initial Neural Network: LLMatic starts off with a simple neural network with one convolutional layer that takes in $3$ input with 3 channels, with $1\times 1$ kernel size and $1$ output channel which connects to a dense layer of size $1024$. These hidden neurons are connected via another dense layer to $10$ output neurons (as we have 10 classes). Rectified Linear Unit (ReLU) (Nair and Hinton, 2010) is the activation function used in all layers. All of our networks are generated in PyTorch (Paszke et al., 2019).
@@
For evolutionary operators, we set a probability of $0.7$ for mutation and $0.3$ for crossover as after experimentation, we found that mutation creates consistently more trainable neural networks. We initialize the temperature parameter (used when sampling the code-generating LLM) to $0.6$. For temperature mutation, half of the population is generated by the prompt individual temperature mutated uniformly at random between $-0.1$ to $0.1$. The other half is generated by the temperature obtained from the prompt individual itself. If the fitness of the generated network is better than or equal to the best fitness of the previous generation, we increase the temperature by $0.05$ and if it is worse than the best fitness of the previous generation, we decrease it by $0.05$. For the crossover operator, we select $10$ random networks and find their $2$ or $3$ nearest neighbours in the network archive to perform crossover. We set the temperature to be $0.7$ for network generation.
Quality Diversity Optimization: For our QD optimization algorithm, we choose a variant of MAP-Elites, called Centroidal Voronoi Tessellation (CVT-MAP-Elites) (Vassiliades et al., 2017) can be seen as a generalization of MAP-Elites, intended to ease scaling up MAP-Elites to high-dimensional behavior spaces; CVT-MAP-Elites outperforms standard MAP-Elites in high-dimensional spaces while matching performance lower-dimensional scenarios such as those studied in the present work. CVT-MAP-Elites automates the sub-division of the archive by identifying $k$ cell centroid locations exhibiting an even spread through the behavioural descriptors space. (Here, $k$ corresponds to the total number of evolutionary “niches” in the archive.) We use the pymap_elites [^1] https://github.com/resibots/pymap_elites implementation for our experimentation. k-d tree (Bentley, 1975) is used to create and write centroids to the archive and find the nearest neighbors using a Euclidean distance metric (Dokmanic et al., 2015).
Each of our QD archives has $2$ dimensions (behavioral descriptors), with $100$ niches spread across them. We set our number of random initial networks to $10$. For the network archive, we have the width-to-depth ratio of the network as our first dimension and the FLOPS of the network as the second dimension. The width-to-depth ratio has a lower limit of 00 and an upper limit of $200$. The minimum FLOPS is set to $200$ MegaFLOPS and the maximum is set to $5$ GigaFLOPS. This range is set after experimentation.
@@
For the network archive, we simply select a random network while for the prompt archive, we select the most curious prompt individual, which depends on the curiosity score. This curiosity score is incremented by $1.0$ if the selected prompt adds the generated network to the network archive, decreased by $0.5$ if the network is not added, and reduced by $1.0$ if the created network is untrainable. If the generated network has better fitness than the previous generation’s best network, the collective prompt fitness score for the prompt in the prompt individual is increased by $1$, otherwise, it is unchanged. We use prompts that are generalizable to any problem in any domain. Refer to LABEL:appendixprompts for an example of mutation and crossover prompts.
Figure 3. The illustration of how many trainable networks are created in a generation. The total number of networks created is 100 per generation.
Code Generating LLM: We use the pre-trained CodeGen (Nijkamp et al., 2022) LLM to generate neural networks. CodeGen is an autoregressive decoder-only transformer with left-to-right causal masking. CodeGen is first trained on ThePile dataset with random initialization and is called CodeGen-NL. CodeGen-Multi is initialized with CodeGen-NL and is trained on BigQuery dataset. Lastly, CodeGen-Mono is initialized with CodeGen-Multi and is trained on BigPython. CodeGen is trained to be in various parameter sizes but we use $6.1$ Billion parameter variant of CodeGen-Mono due to computational constraints.
ThePile dataset (Gao et al., 2020) is an $825.18$ GB English text corpus. (Nijkamp et al., 2022) select a subset of the Google BigQuery dataset which contains 6 programming languages, namely C, C++, Go, Java, JavaScript, and Python. The authors collected a large amount of permissively licensed Python code from GitHub in October $2021$, and named it BigPython. The size of BigPython is $217.3$ GB.
@@ 4.2. Ablation Study
  • Network-Archive-LLMatic: LLMatic with only the network archive. To achieve this, we create a population of prompt individuals. The population is fixed to $100$ individuals initialized with random individuals. We have only one fitness score for this population, which is calculated as $+1$ if a network is added in the network archive, $-0.5$ if the network is not added and $-1$ if the network is not trainable. After we generate the network, we mutate temperature by adding $0.1$ if the network is added in the network archive and $-0.1$ if the network is not added.
@@
4.2.1. Ablation Results and Discussion
In this section, we will discuss the results of the experiments that we set up in the previous section. We first discuss the best loss per generation, illustrated in Figure 2. This will lead our discussion to trainable networks generated by changing the crossover and mutation probabilities (Figure 3). Then we will discuss how archives are illuminated Figure 4. Some of the generated networks are shown in the supplementary material.
@@ 4.2.1. Ablation Results and Discussion
Figure 4. An illustration of archives generated by LLMatic. We have selected the archive with the median number of cells filled in experiments over 30 seeds. 3(a) shows prompt archive. 3(b) shows network archive. The lighter the colour of the filled cell, the better fitness of the individual.
@@
Looking at Figure 2, it is clear that each component of LLMatic is necessary. Mutation-Only-LLMatic and Network-Archive-LLMatic are the closest to LLMatic which also indicates that our choice of giving mutation more probability of being selected is the right one. Crossover-Only-LLMatic is understandably the worse as mutation provides more exploration (Ullah et al., 2022). Both operators, mutation and crossover, together give exploration and exploitation abilities to LLMatic, which are highly necessary to find high-quality and diverse networks. While Prompt-Archive-LLMatic does significantly worse as network archive is an important aspect to find high-performing networks. Both archives together demonstrate competitive results.
@@
We use EfficientNet-B0, which is the state-of-the-art network on CIFAR-10 Tan and Le (2019) as an indicator of where our algorithm stands. EfficientNet-B0 was searched via methods applied by Tan et al. (2019) and is slightly larger than the original study as they were targeting more FLOPS. The original study required $8,000$, while LLMatic requires $2,000$ evaluations to find a competitive network. EfficientNet-B0 was first trained on the ImageNet dataset (Deng et al., 2009) and then on CIFAR-10 via transfer learning (Torrey and Shavlik, 2010). This is an advantage for EfficientNet-B0 as ImageNet has many classes and is an order of magnitude larger dataset.
Figure 4 demonstrates how each archive is filled on average. We can see the prompt archive contains high-performing individuals who have the first few prompts and higher temperatures.
@@
Some of the high-performing individuals do have lower temperatures, which suggests that sometimes it is good to pick deterministic layers. For network archives, we observe a diversity of high-performing networks with respect to both FLOPS and width-to-depth ratio. More than $20$ networks are competitive networks in this archive.
@@
To delve into why we choose the probabilities being $0.7$ for mutation and $0.3$ for crossover, we observe the number of trainable networks generated per generation (see Figure 3). This is to be considered common knowledge that the more working individuals we have, the greater the chance of high-performing individuals. For this purpose, we train LLMatic with uniform probabilities, and $0.3$ for mutation and $0.7$ for crossover. We observe that uniform probabilities are still competitive with the original setting, while increasing the crossover probability makes it worse. The results of these experiments and results of the ablation study for Mutation-Only-LLMatic and Crossover-Only-LLMatic lead us to the conclusion that mutation should be given more probability of being selected.

### v8

@@ 1. Introduction
One could argue that search by evolutionary computation or reinforcement learning is quite mindless and wasteful, given how many architectures need to be tested and how uninformed the changes that lead to each new architecture are. Is there some way we can inform the search by exploiting stored knowledge about how to design neural networks? This paper explores the idea that we can do exactly this using code-generating large language models (LLMs). More precisely, we propose combining an LLM with an evolutionary algorithm to generate new architectures that have high network architectural diversity and state-of-the-art performance.
@@ 2. Related Work
To avoid having the designer constrain the space of possible architectures prior to search, we turn to code-generating large language models (LLMs)—large models trained auto-regressively on massive datasets of code (e.g. public repositories hosted on Github). These LLMs are based on the transformer architecture (Vaswani et al., 2017) that has obtained state-of-the-art performance in natural language modelling (Radford et al., 2019; Brown et al., 2020; Adelani et al., 2022; Nasir and Mchechesi, 2022). They have also been successfully used in specific applications, such as for video game level design (Todd et al., 2023; Nasir and Togelius, 2023; Chen et al., 2021) or code generation (Shinn et al., 2023).
@@ 3. Approach
LLMatic begins its search with a very basic neural network, inspired by the work of  (Stanley and Miikkulainen, 2002) which suggests that neuroevolution tends to perform better when starting with a small network. In LLMatic, we use a novel dual-archive cooperative QD optimization approach, in which two separate archives are used to store complementary components that can be combined to solve a given task. The first archive stores neural networks, where the width-to-depth ratio and Floating Point Operations per Second (FLOPS) of a network are the behavioural descriptors. The width-to-depth ratio is a division of the width and the depth of the network. To specify the width, we use the maximum of the output features of all layers, while depth is simply the number of layers. Note that we choose FLOPS instead of parameter count because FLOPS correlates better with actual time spent training a network (Ayala et al., 2017). We call this archive the “network archive”. The fitness function for the networks in this archive is defined as the test accuracy of the network after training. The second archive, called the “prompt archive”, contains the prompt and temperature used for generating code, which are also the behavioural descriptors. The temperature of an LLM is a hyperparameter that controls the degree of stochasticity in the selection of output tokens: a lower temperature means that the LLM will select tokens more deterministically, while a higher value will result in more diverse output. The selection of prompt and temperature depends on a curiosity score (Cully and Demiris, 2017), governed by whether the generated network was added to the network archive. The fitness of prompt individuals depends on whether the network was better than the previous generation’s best score. Figure 1 illustrates the flowchart of the approach LLMatic uses, while Algorithm 1 shows the complete search process of LLMatic in pseudocode.
In the first generation, a simple neural network with one convolutional and one fully connected layer initiates the evolution (line 1 of Algorithm 1). A prompt is selected at random to generate an initial batch of networks (lines 5–6 of Algorithm 1). These networks are evaluated and an attempt is made to add them to the network archive as a random initialization for MAP-Elites. Concurrently, we mutate the temperature based on the fitness of the network, increasing it if the fitness increases and vice versa (lines 21–23 of Algorithm 1). An increase in temperature is desirable if we want the LLM to explore, while decreasing the temperature will result in the LLM exploiting in an attempt to achieve better fitness than before. Once we calculate the fitness of the prompt individual, we add the score to a collective prompt fitness score, after which we try to populate the prompt archive. The collective prompt fitness score determines the overall fitness of each individual in the prompt archive as it gives each prompt a fitness score.
Once either of the archives reaches a specified capacity, we introduce neural network training and evolutionary operators in the process (lines 7–20 of Algorithm 1). With a certain probability at each generation, a decision is made on whether to perform crossover or mutation to produce $N$ new offspring. If the crossover operator is chosen, we select $N$ random network individuals, locate their closest networks in the archive, and carry out a crossover operation instructed by a prompt (lines 15–16 of Algorithm 1). No individual is added to the prompt archive when a crossover is performed. If the mutation operation is selected, we pick the most curious prompt individual and a random network individual. For exploration, we also select random prompts. In both cases, each network is trained for a certain number of epochs and an attempt is made to add the network to the archive. Likewise, a prompt individual is added as previously described. This process continues for a predetermined number of generations. Refer to the supplementary material for pseudocode on mutation operators, crossover operators, temperature mutation and addition to archives.
1   Initialize network and prompt archives.
2while number of generations $<$ maximum generations
3
@@
7
8                  Mutate the initial network with a random prompt.
9                  Get the network and prompt individuals and add them to the respective archives.
@@
11                   Randomly choose mutation or crossover as the evolutionary operator.
12                  if evolutionary operator ==absent$=$ = mutation then
13
14                        Select network and prompt.
15                        Mutate the selected network with the selected prompt.
16                        Train or query the generated network.
17                        Get the network and prompt individuals and store them.
@@
19
20                        Selection of networks for crossover.
21                        Perform crossover on selected networks with a fixed prompt.
22                        Train or query the generated network.
23                        Store the generated network individual.
@@
29
30      Evaluate all networks to find the test accuracies.
31      Mutate the temperature for the next generation.
32      Add the batch network individuals and the corresponding prompt individuals in the respective archives.
@@ 4.1. Setting up LLMatic
Initial Neural Network: LLMatic starts off with a simple neural network with one convolutional layer that takes in $3$ input channels, with $1\times 1$ kernel size and $1$ output channel which connects to a dense layer of size $1024$. These hidden neurons are connected via another dense layer to $10$ output neurons (as we have 10 classes). Rectified Linear Unit (ReLU) (Nair and Hinton, 2010) is the activation function used in all layers. All of our networks are generated in PyTorch (Paszke et al., 2019).
@@
For evolutionary operators, we set a probability of $0.7$ for mutation and $0.3$ for crossover as after experimentation, we found that mutation creates consistently more trainable neural networks. We initialize the temperature parameter (used when sampling the code-generating LLM) to $0.6$. For temperature mutation, half of the population is generated by the prompt individual temperature mutated uniformly at random between $-0.1$ to $0.1$. The other half is generated by the temperature obtained from the prompt individual itself. If the fitness of the generated network is better than or equal to the best fitness of the previous generation, we increase the temperature by $0.05$ and if it is worse than the best fitness of the previous generation, we decrease it by $0.05$. For the crossover operator, we select $10$ random networks and find their $2$ or $3$ nearest neighbours, based on the distance of niches, in the network archive to perform crossover. We set the LLM temperature to be $0.7$ for network generation.
Quality Diversity Optimization: For our QD optimization algorithm, we choose a variant of MAP-Elites—Centroidal Voronoi Tessellation (CVT-MAP-Elites) (Vassiliades et al., 2017)—which can be seen as a generalization of MAP-Elites that is intended to scale MAP-Elites to high-dimensional behavior spaces; CVT-MAP-Elites outperforms standard MAP-Elites in high-dimensional spaces while matching performance lower-dimensional scenarios such as those studied in the present work.
CVT-MAP-Elites automates the sub-division of the archive by identifying $k$ cell centroid locations exhibiting an even spread through the behavioural descriptors space. Here, $k$ corresponds to the total number of evolutionary “niches” in the archive. We use the pymap_elites [^1] https://github.com/resibots/pymap_elites implementation for our experimentation. We use a k-d tree (Bentley, 1975) to create and write centroids to the archive and find the nearest neighbors using a Euclidean distance metric (Dokmanic et al., 2015).
Each of our QD archives has $2$ dimensions (behavioral descriptors), with $100$ niches spread across them. We set the number of random initial networks to $10$. For the network archive, we have the width-to-depth ratio of the network as our first dimension and the FLOPS of the network as the second dimension. The width-to-depth ratio has a lower limit of 00 and an upper limit of $200$. The minimum FLOPS is set to $200$ MegaFLOPS and the maximum is set to $5$ GigaFLOPS. This range is set after experimentation.
@@
For the network archive, we simply select a random network while for the prompt archive, we select the most curious prompt individual, which depends on the curiosity score. This curiosity score is incremented by $1.0$ if the selected prompt adds the generated network to the network archive, decreased by $0.5$ if the network is not added, and reduced by $1.0$ if the created network is untrainable. If the generated network has better fitness than the previous generation’s best network, the collective prompt fitness score for the prompt in the prompt individual is increased by $1$; otherwise, it is unchanged. We use prompts that are generalizable to any problem in any domain. Refer to supplementary material for an example of mutation and crossover prompts.
Code Generating LLM: We use the pre-trained CodeGen (Nijkamp et al., 2022) LLM to generate neural networks. CodeGen is an autoregressive decoder-only transformer with left-to-right causal masking. CodeGen is first trained on ThePile dataset with random initialization and is called CodeGen-NL. CodeGen-Multi is initialized with CodeGen-NL and is trained on BigQuery dataset. Lastly, CodeGen-Mono is initialized with CodeGen-Multi and is trained on BigPython. CodeGen is trained to be in various parameter sizes, but we use $6.1$ Billion parameter variant of CodeGen-Mono due to computational constraints.
ThePile dataset (Gao et al., 2020) is an $825.18$ GB English text corpus. CodeGen selects a subset of the Google BigQuery dataset which contains 6 programming languages, namely C, C++, Go, Java, JavaScript, and Python. The authors collected a large amount of permissively licensed Python code from GitHub in October $2021$, and named it BigPython. The size of BigPython is $217.3$ GB.
@@ 4.2. Ablation Study
  • Network-Archive-LLMatic: LLMatic with only the network archive. To achieve this, we create a population of prompt individuals. The population is fixed to $100$ individuals initialized with random individuals. We have only one fitness score for this population, which is calculated as $+1$ if a network is added in the network archive, $-0.5$ if the network is not added and $-1$ if the network is not trainable. After we generate the network, we mutate the temperature by adding $0.1$ if the network is added in the network archive and $-0.1$ if the network is not added.
@@
Figure 3. An illustration of archives generated by LLMatic. We have selected the archive with the median number of cells filled in experiments over 30 seeds. 3(a) shows the prompt archive, while 3(b) shows the network archive. The lighter the colour of the filled cell, the better fitness of the individual. White indicates that the cell is empty.
4.2.1. Ablation Results and Discussion
In this section, we will discuss the results of the experiments that we set up in the previous section. We first discuss the best accuracy per generation, illustrated in Figure 2. This will lead our discussion to trainable networks generated by changing the crossover and mutation probabilities (Figure 4). Then we will discuss how archives are illuminated Figure 3. Some of the generated networks are shown in the supplementary material.
@@ 4.2.1. Ablation Results and Discussion
Figure 2 illustrates that each component of LLMatic is necessary. Mutation-Only-LLMatic and Network-Archive-LLMatic are the closest to LLMatic, which validates our choice to weight the probability of mutation higher. Crossover-Only-LLMatic performs the worst, as it does not benefit from the exploration abilities provided by the mutation operator (Ullah et al., 2022). Both operators (mutation and crossover) together provide exploration and exploitation abilities to LLMatic, which appear necessary to find high-quality and diverse networks. Prompt-Archive-LLMatic performs poorly, indicating that the network archive is an important aspect in finding high-performing networks. However, both archives together demonstrate competitive results.
@@
We use EfficientNet-B0, which is the state-of-the-art network on CIFAR-10 Tan and Le (2019) as an indicator of where our algorithm stands. EfficientNet-B0 was searched via methods applied by Tan et al. (2019) and is slightly larger than the original study as they were targeting more FLOPS. The original study required $8,000$ evaluations, while LLMatic requires $2,000$ evaluations to find a competitive network. EfficientNet-B0 was first trained on the ImageNet dataset (Deng et al., 2009) and then on CIFAR-10 via transfer learning (Torrey and Shavlik, 2010). This is an advantage for EfficientNet-B0 as ImageNet has many classes and is an order of magnitude larger dataset.
@@
Figure 3 demonstrates how each archive is filled on average. We can see that the prompt archive contains high-performing individuals who have the first few prompts and higher temperatures. Some of the high-performing individuals do have lower temperatures, which suggests that sometimes it is useful to generate neural network layers in a less stochastic manner. For network archives, we observe a diversity of high-performing networks with respect to both FLOPS and width-to-depth ratio. More than $20$ individuals are competitive networks in this archive.
@@
To investigate our choice of probabilities for crossover and mutation ($0.3$ and $0.7$, respectively), we observe the number of trainable networks generated per generation (see Figure 4). We use this as a measure, since the more functional individuals we have, the greater the chance of high-performing individuals. For this purpose, we train LLMatic with uniform probabilities, and $0.3$ for mutation and $0.7$ for crossover. We observe that uniform probabilities are still competitive with the original setting, while increasing the crossover probability makes it worse. The results of these experiments and results of the ablation study for Crossover-Only-LLMatic and Mutation-Only-LLMatic lead us to the conclusion that mutation should be given more probability of being selected.
Figure 4. The illustration of how many trainable networks are created in a generation. The total number of networks created is 100 per generation. This illustration is calculated over 10 runs. The shaded region is the standard deviation.